1. how long does it take 1700 to double if it invested at 5% interest , compound monthly?
2. what annual rate of interest is required to doulbe an investment in 7 years?

Answer by lwsshak3(11628)

1. how long does it take 1700 to double if it invested at 5% interest , compound monthly?
2. what annual rate of interest is required to double an investment in 7 years?
..
Compound interest formula:
A=P(1+i)^n,A= amount of investment after n periods, P=original investment, i=interest rate per period, n=no of periods.
..
interest/period=.05/12
no of periods = n(months)
(1+i)=(1+.05/12)=1.00417
1) A/P=(1+i)^n
2=1.00417^n
log2=nlog1.00417
n=log2/log1.00417
n=166.56 months=166/12=13.88 years
ans: It will take around 13.9 years to double the 1700 investment at the annual interest rate of 5%
..
2) 7 years=84 months=n
2=(1+i)^84
raise both sides by 1/84
2^(1/84)=(1+i)^84)^(1/84)
1.00829=1+i
i=.00829 =monthly interest
annual interest=12*.00829=.099=9.9%
Ans:
Annual interest required to double investment in 7 years is about 10%